# Yuchen Zhang

### v2

@@ 4.2 Artifact Pre-perception Encoding
This stage aims to perceive manipulation artifacts within input data and condense these forensic clues into the artifact token. Specifically, the input sequence $S$ undergoes processing through the artifact pre-perception multimodal encoder $\mathcal{E}^{p}_{m}$, yielding $\hat{S}=[\hat{E}_{v};\hat{E}_{a};\hat{E}_{t}]$. To inject artifact-aware clues into the artifact token embedding $\hat{E}_{a}$, we pick $\hat{E}_{a}$ from $\hat{S}$ and feed it into an artifact-aware classification head. As illustrated in Fig. 3, this classification head is optimized via a manipulation detection objective to explicitly encode artifact-related patterns into $\hat{E}_{a}$.
@@ 4.3 Manipulation-Oriented Reasoning
Manipulation-Oriented Reasoning (MOR) is in charge of generating the textual answer in response to the question prompt. To acquire an accurate response, we augment the network optimization in MOR with two guiding tasks: visual Artifact Capture via Grounding and Manipulation-focused Guidance.
@@
Subsequently, the $u_{agg}$ is sent to the bbox detector to generate artifact coordinates. We follow [Loss_IoU] to construct the image manipulation grounding loss using L1 loss $\mathcal{L}_{1}$ and GIoU loss $\mathcal{L}_{IoU}$:
@@
$u_{v}=\text{Attention}(\hat{E}_{v+a}^{m},\hat{E}_{t}^{m},\hat{E}_{t}^{m}),\quad u_{t}=\text{Attention}(\hat{E}_{t}^{m},\hat{E}_{v+a}^{m},\hat{E}_{v+a}^{m}),$ (5)
@@
Language modeling. The input sequence $S_{a}$ is processed through an encoder-decoder architecture, ultimately generating a pure text output that includes choices and coordinates (Fig. 3d) as specified in the prompts. In this stage, an autoregressive approach is adopted, where the decoder generates the target sequence $y$ conditioned on $S_{m}$. The language modeling loss $\mathcal{L}_{LM}$ [Florence2_CVPR] is used to supervise the decoded text outputs.
@@ 4.4 Token Redundancy Penalty
To suppress the redundancy and increase the information density among tokens in $E_{a}$, we design a Token Redundancy Penalty (TRP) optimization term. Specifically, we first encourage the columns of $E_{a}$ to be as orthogonal as possible by introducing a loss term $\mathcal{L}_{\mathrm{orth}}$, which increases the matrix rank. We construct Gram matrix of $E_{a}$, $G=E_{a}E_{a}^{\top}\in\mathbb{R}^{n_{a}\times n_{a}}$, and the orthogonality of the columns can be measured by the off-diagonal elements of the Gram matrix. Ideally, if the columns are orthogonal, the off-diagonal entries of $G$ should be zero. Therefore, we define:
@@
$\mathcal{L}_{\mathrm{mod}}=\frac{1}{n_{a}}\sum_{t=1}^{n_{a}}\left(\sum_{i=1}^{d}p_{t,i}\log p_{t,i}+\log d\right),$ (8)
@@
$\mathcal{L}_{{TRP}}=\mathcal{L}_{\mathrm{orth}}+\mathcal{L}_{\mathrm{mod}},$ (9)
$\mathcal{L}_{TRP}$ is imposed on the $S_{inp}$ sequence during the APE stage.
4.5 Training and Inference
@@
$\mathcal{L}=\mathcal{L}_{APE}+\mathcal{L}_{DBM}+\mathcal{L}_{IMG}+\mathcal{L}_{TRP}+\mathcal{L}_{LM},$ (10)
@@ 5 Experiment
Zero-Shot Qwen2.5-VL  [bai2025qwen25] 47.74 29.24 0.00 35.18 25.70 0.00 24.66 40.60 0.00 25.11 40.29 0.28 35.89 31.51 0.00 33.72 33.47 0.06
Qwen3-a22b  [Qwen3_blog] 45.29 25.01 0.71 38.12 27.41 1.19 22.87 39.17 0.10 22.87 40.77 1.34 37.22 31.60 0.98 33.27 33.69 0.86
GPT-4o  [hurst2024gpt4o] 48.48 27.90 0.82 35.68 29.49 1.23 24.62 39.88 1.37 23.62 38.89 1.22 37.19 30.48 1.20 33.92 33.33 1.17
Gemini-2.0  [team2023gemini] 56.05 33.16 1.44 41.26 24.37 1.12 29.60 38.29 1.40 29.15 35.20 2.42 38.12 29.13 2.25 38.83 32.03 1.72
Tr. on NYT ViLT  [ViLT_ICML] 83.27 64.27 22.73 72.18 31.76 20.21 70.34 36.45 21.48 65.71 36.23 17.56 74.33 36.10 19.36 73.17 40.96 20.27
HAMMER  [DGM4] 79.20 55.86 51.34 68.23 40.10 21.56 71.52 41.17 13.74 68.50 41.47 13.92 67.37 42.23 16.12 70.96 44.16 23.34
HAMMER++  [DGM4_TPAMI] 79.61 57.06 54.44 66.99 38.07 17.34 67.18 37.58 10.76 66.28 37.97 10.88 66.12 37.82 13.68 69.23 41.70 21.42
FKA-Owl  [liu2024fkaowl] 94.67 78.18 55.81 77.20 46.88 43.67 78.00 44.45 50.73 75.49 50.83 43.53 84.65 60.73 43.28 81.60 56.77 46.23
@@
Tr. on Guardian ViLT  [ViLT_ICML] 68.80 43.99 21.77 85.29 67.34 41.80 70.34 46.24 37.68 78.61 47.17 38.13 80.00 44.79 38.97 76.61 49.90 35.67
HAMMER  [DGM4] 61.89 37.98 18.84 78.50 52.40 51.53 74.78 50.76 43.40 75.11 50.34 46.36 81.32 50.15 56.03 74.32 48.33 43.23
HAMMER++  [DGM4_TPAMI] 62.75 36.45 23.76 80.95 59.92 64.67 75.36 48.77 47.13 76.30 49.56 48.91 80.12 50.36 57.97 75.10 49.01 48.49
FKA-Owl  [liu2024fkaowl] 80.60 40.44 26.33 92.60 78.24 71.04 80.90 51.80 50.93 78.88 51.62 50.88 87.61 68.57 61.82 84.12 58.13 52.20
@@
ViLT  [ViLT_ICML] 68.27 42.29 69.87 43.19 52.79 31.28 62.11 33.78 55.76 33.26 57.17 31.10 44.14 39.68 59.06 21.96 55.24 36.63 62.05 32.49
HAMMER  [DGM4] 78.34 66.79 78.27 61.09 64.97 40.49 73.76 40.51 63.54 40.26 76.13 38.53 54.97 40.84 81.48 43.74 65.45 47.10 77.41 45.97
HAMMER++  [DGM4_TPAMI] 79.13 67.11 78.24 62.15 65.25 40.74 73.24 41.14 63.83 40.34 76.17 38.21 54.24 41.25 81.73 43.23 65.61 47.36 77.34 46.19
FKA-Owl  [liu2024fkaowl] 82.97 53.86 87.70 65.69 67.57 38.97 79.44 32.57 67.05 37.70 81.55 31.86 70.26 40.20 84.54 46.48 71.96 42.68 83.31 44.15
@@
Table 4: Ablation on (a) each components and (b) discussion regarding performance on test set of difference MLLMs & differnece manipulation type.
@@
Please refer to the appendix for the experimental setup and evaluation metrics.
@@ 5.1 Quantitative Results
Effectiveness & Generalization. We assess AMD against four SOTA methods on the MDSM and $\text{DGM}^{4}$ datasets. For MDSM (Tab. 2), we train on The Guardian and NYT, testing on the rest. For $\text{DGM}^{4}$ (Tab. 3), we train on the largest subset, The Guardian. Tab. 2 also shows zero-shot results for general-purpose models. Our key findings are: (1) MLLMs’ knowledge boosts performance. Forgery-trace methods like ViLT [ViLT_ICML] and HAMMER series [DGM4_TPAMI] show limited performance, unlike MLLM-based methods like FKA-Owl [liu2024fkaowl] and AMD. For instance, trained on MDSM-NYT (Tab. 2), AMD achieves an 84.47 intra-domain mAP and >60 cross-domain, while HAMMER scores 57.06 and <42, respectively. (2) AMD achieves strong grounding. AMD attains the best average mIoU of 63.56 (NYT-trained) and 61.02 (Guardian-trained) (Tab. 2). General-purpose models perform poorly (mIoU < 3). AMD’s superiority stems from its question-answer heuristic prompts and MOR module, which omits coordinate outputs when no manipulation is detected, thus reducing unnecessary errors. (3) AMD generalizes effectively. On $\text{DGM}^{4}$ (Tab. 3), AMD outperforms the HAMMER series on all metrics (74.47 ACC, 52.91 mAP, 80.01 $\text{P}_{tok}$, 51.87 mIoU). It also surpasses FKA-Owl in ACC, mAP, and mIoU, despite a lower $\text{P}_{tok}$.
Generalization Assessment across MLLMs. To assess generalization on different MLLMs, we evaluated an NYT-trained AMD on test narratives generated by four MLLMs: Qwen-VL [QwenVL], X-InstructBLIP [xInstructBLIP_2024], LLaVA [LLAVA_NeurIPS], and mPLUG-Owl [ye2024mplugowl3longimagesequenceunderstanding]. Results (Tab. 4b, chart I) show robust performance, with intra-domain (NYT) and cross-domain APs exceeding 76 and 53, respectively.
@@
5.2 Ablation Study
Figure 4: Visualization of Artifact Token sequence. (a) Visualization of $E_{a}$ in $S_{inp}$. (b) Visualization of $\hat{E_{a}}$ in $S_{a}$. (c) Visualization of $\hat{E_{a}^{m}}$ in $S_{m}$
@@ 5.2 Ablation Study
Artifact Token Visualization. Fig. 4 visualizes the Artifact token at different stages via t-SNE [tsne_JMLR]. As shown in Fig. 4a to c, the sample points progressively form more distinct clusters, clearly demonstrating the effectiveness of our AMD optimization in enhancing the Artifact Token’s ability to distinguish between different categories.
For a more complete understanding of the MDSM dataset and the ablations on the AMD design, we refer readers to the appendix.

### v4

@@ 3.2 Artifact Pre-perception Encoding
This stage aims to perceive manipulation artifacts within input data and condense these forensic clues into the artifact token. Specifically, the input sequence $S$ undergoes processing through the artifact pre-perception multimodal encoder $\mathcal{E}^{p}_{m}$, yielding $\hat{S}=[\hat{E}_{v};\hat{E}_{a};\hat{E}_{t}]$. To inject artifact-aware clues into the artifact token embedding $\hat{E}_{a}$, we pick $\hat{E}_{a}$ from $\hat{S}$ and feed it into an artifact-aware classification head. As illustrated in Fig. 4, this classification head is optimized via a manipulation detection objective to explicitly encode artifact-related patterns into $\hat{E}_{a}$.Then
@@ 3.3 Manipulation-Oriented Reasoning
Manipulation-Oriented Reasoning (MOR) targets to generate the textual answer in response to the question prompt. To acquire an accurate response, we augment the network optimization in MOR with two guiding tasks: visual Artifact Capture via Grounding and Manipulation-focused Guidance.
@@
Subsequently, the $u_{agg}$ is sent to the bbox detector to generate artifact coordinates. We follow [36] to construct the image manipulation grounding loss using L1 loss $\mathcal{L}_{1}$ and GIoU loss $\mathcal{L}_{IoU}$:
@@
$u_{v}$ $=\text{Attention}(\hat{E}_{v+a}^{m},\hat{E}_{t}^{m},\hat{E}_{t}^{m}),$ (5)
$u_{t}$ $=\text{Attention}(\hat{E}_{t}^{m},\hat{E}_{v+a}^{m},\hat{E}_{v+a}^{m}),$
@@
Language modeling. The input sequence $S_{a}$ is processed through an encoder-decoder architecture, ultimately generating a pure text output that includes choices and coordinates (Fig. 4d) as specified in the prompts. In this stage, an autoregressive approach is adopted, where the decoder generates the target sequence $y$ conditioned on $S_{m}$. The language modeling loss $\mathcal{L}_{LM}$ [5] is used to supervise the text outputs.
@@ 3.4 Token Redundancy Penalty
To suppress the redundancy and increase the information density among tokens in $E_{a}$, we design a Token Redundancy Penalty (TRP) optimization term. Specifically, we first encourage the columns of $E_{a}$ to be as orthogonal as possible by introducing a loss term $\mathcal{L}_{\mathrm{orth}}$, which increases the matrix rank. We construct Gram matrix of $E_{a}$, $G=E_{a}E_{a}^{\top}\in\mathbb{R}^{n_{a}\times n_{a}}$, and the orthogonality of the columns can be measured by the off-diagonal elements of the Gram matrix. Ideally, if the columns are orthogonal, the off-diagonal entries of $G$ should be zero. To this end, we define following constraint:
@@
$\mathcal{L}_{\mathrm{mod}}=\frac{1}{n_{a}}\sum_{t=1}^{n_{a}}\left(\sum_{i=1}^{d}p_{t,i}\log p_{t,i}+\log d\right).$ (8)
@@
$\mathcal{L}_{{TRP}}=\mathcal{L}_{\mathrm{orth}}+\mathcal{L}_{\mathrm{mod}}.$ (9)
@@
$\mathcal{L}=\mathcal{L}_{APE}+\mathcal{L}_{DBM}+\mathcal{L}_{IMG}+\mathcal{L}_{TRP}+\mathcal{L}_{LM}.$ (10)
@@ 4 Experiment
Zero-Shot LLaVA-v1.6-13B  [22] 32.10 21.86 0.00 21.76 19.18 0.00 17.80 12.76 0.00 15.76 25.47 0.00 22.70 17.38 0.00 22.02 19.33 0.00
DeepSeek-VL2-27B  [54] 38.12 21.79 0.45 30.35 20.12 0.76 22.49 37.27 0.43 21.73 20.02 1.78 29.70 26.58 1.12 28.48 25.16 0.91
Yi-VL-34B  [1] 45.74 24.19 0.07 34.98 23.98 0.13 20.63 36.74 0.00 19.28 19.10 0.00 31.39 26.97 0.00 30.40 26.20 0.04
Qwen2.5-VL-72B  [4] 47.74 29.24 0.00 35.18 25.70 0.00 24.66 40.60 0.00 25.11 40.29 0.28 35.89 31.51 0.00 33.72 33.47 0.06
Qwen3-VL-235B  [55] 45.29 25.01 0.71 38.12 27.41 1.19 22.87 39.17 0.10 22.87 40.77 1.34 37.22 31.60 0.98 33.27 33.69 0.86
GPT-4o  [16] 48.48 27.90 0.82 35.68 29.49 1.23 24.62 39.88 1.37 23.62 38.89 1.22 37.19 30.48 1.20 33.92 33.33 1.17
Gemini-2.0  [46] 56.05 33.16 1.44 41.26 24.37 1.12 29.60 38.29 1.40 29.15 35.20 2.42 38.12 29.13 2.25 38.83 32.03 1.72
Tr. on NYT ViLT  [18] 83.27 64.27 22.73 72.18 31.76 20.21 70.34 36.45 21.48 65.71 36.23 17.56 74.33 36.10 19.36 73.17 40.96 20.27
HAMMER  [40] 79.20 55.86 51.34 68.23 40.10 21.56 71.52 41.17 13.74 68.50 41.47 13.92 67.37 42.23 16.12 70.96 44.16 23.34
HAMMER++  [41] 79.61 57.06 54.44 66.99 38.07 17.34 67.18 37.58 10.76 66.28 37.97 10.88 66.12 37.82 13.68 69.23 41.70 21.42
FKA-Owl  [23] 94.67 78.18 55.81 77.20 46.88 43.67 78.00 44.45 50.73 75.49 50.83 43.53 84.65 60.73 43.28 81.60 56.77 46.23
@@
Tr. on Guardian ViLT  [18] 68.80 43.99 21.77 85.29 67.34 41.80 70.34 46.24 37.68 78.61 47.17 38.13 80.00 44.79 38.97 76.61 49.90 35.67
HAMMER  [40] 61.89 37.98 18.84 78.50 52.40 51.53 74.78 50.76 43.40 75.11 50.34 46.36 81.32 50.15 56.03 74.32 48.33 43.23
HAMMER++  [41] 62.75 36.45 23.76 80.95 59.92 64.67 75.36 48.77 47.13 76.30 49.56 48.91 80.12 50.36 57.97 75.10 49.01 48.49
FKA-Owl  [23] 80.60 40.44 26.33 92.60 78.24 71.04 80.90 51.80 50.93 78.88 51.62 50.88 87.61 68.57 61.82 84.12 58.13 52.20
@@
ViLT  [18] 68.27 42.29 69.87 43.19 52.79 31.28 62.11 33.78 55.76 33.26 57.17 31.10 44.14 39.68 59.06 21.96 55.24 36.63 62.05 32.49
HAMMER  [40] 78.34 66.79 78.27 61.09 64.97 40.49 73.76 40.51 63.54 40.26 76.13 38.53 54.97 40.84 81.48 43.74 65.45 47.10 77.41 45.97
HAMMER++  [41] 79.13 67.11 78.24 62.15 65.25 40.74 73.24 41.14 63.83 40.34 76.17 38.21 54.24 41.25 81.73 43.23 65.61 47.36 77.34 46.19
FKA-Owl  [23] 82.97 53.86 87.70 65.69 67.57 38.97 79.44 32.57 67.05 37.70 81.55 31.86 70.26 40.20 84.54 46.48 71.96 42.68 83.31 44.15
@@
Table 4: Ablation on (a) each component and (b) discussion regarding performance on test set of different MLLMs & manipulation types.
@@
For the implementation details of all experimental models, the training settings, and the definitions of the evaluation metrics, please refer to the appendix.
@@ 4.1 Quantitative Results
Effectiveness & Generalization. We assess AMD against four SOTA methods on the MDSM and $\text{DGM}^{4}$ datasets. For MDSM (Tab. 2), we train on The Guardian and NYT, testing on the rest. For $\text{DGM}^{4}$ (Tab. 3), we train on the largest subset, The Guardian. Tab. 2 also shows zero-shot results for general-purpose models. Our key findings are: (1) MLLMs’ knowledge boosts performance. Forgery-trace methods like ViLT [18] and HAMMER series [41] show limited performance, unlike MLLM-based methods like FKA-Owl [23] and AMD. For instance, trained on MDSM-NYT (Tab. 2), AMD achieves an 84.47 intra-domain mAP and >60 cross-domain, while HAMMER scores 57.06 and <42, respectively. (2) AMD achieves strong grounding. AMD attains the best average mIoU of 63.56 (NYT-trained) and 61.02 (Guardian-trained) (Tab. 2). General-purpose models perform poorly (mIoU < 3). AMD’s superiority stems from its question-answer heuristic prompts and MOR module, which omits coordinate outputs when no manipulation is detected, thus reducing unnecessary errors. (3) AMD generalizes effectively. On $\text{DGM}^{4}$ (Tab. 3), AMD outperforms the HAMMER series on all metrics (74.47 ACC, 52.91 mAP, 80.01 $\text{P}_{tok}$, 51.87 mIoU). It also surpasses FKA-Owl in ACC, mAP, and mIoU, despite a lower $\text{P}_{tok}$.
Generalization Assessment across MLLMs. To assess generalization on different MLLMs, we evaluated an NYT-trained AMD on test narratives generated by four MLLMs: Qwen-VL [3], X-InstructBLIP [31], LLaVA [22], and mPLUG-Owl [56]. Results (Tab. 4b, chart I) show robust performance, with intra-domain (NYT) and cross-domain APs exceeding 76 and 53, respectively.
@@
4.2 Ablation Studies
Figure 5: Visualization of Artifact Tokens. (a) Visualization of $E_{a}$ in $S_{inp}$. (b) Visualization of $\hat{E_{a}}$ in $S_{a}$. (c) Visualization of $\hat{E_{a}^{m}}$ in $S_{m}$.
@@ 4.2 Ablation Studies
Artifact Token Visualization. Fig. 5 visualizes the Artifact token at different stages via t-SNE [47]. As shown in Fig. 5a to c, the sample points progressively form more distinct clusters, clearly demonstrating the effectiveness of our AMD optimization in enhancing the Artifact Token’s ability to distinguish between different categories.
Efficiency Discussion. Tab. 5 compares params scale and throughput (images-text pairs processed per second) on RTX 4090. With 276M parameters, AMD is smaller than FKA-Owl (6,771M), enabling faster training and inference. Among comparable-sized models like ViLT and HAMMER, AMD achieves slower speed than HAMMER but significantly outperforms them on misinformation detection tasks. Overall, AMD delivers strong detection performance while maintaining a compact architecture and efficient inference.
Table 5: Efficiency comparison of multi-modal learning methods.
Method Params (M) $\downarrow$ Throughput (p/s) $\uparrow$
Total Trainable Train Inference
ViLT 121.07 121.07 1.85 2.38
HAMMER(++)1 441.12 228.25 28.97 61.28
FKA-Owl 6771.98 33.55 1.25 1.33
AMD 276.95 276.95 5.55 13.38